OpenStudy (anonymous):

What is the solution to the equation log 2 x - log 2 4 = 5 ?

OpenStudy (fifciol):

what is \[\log_{2} 4\] ?

OpenStudy (fifciol):

well it's \[\log_{2} 2^2=2\log_22\]

OpenStudy (fifciol):

\[2\log_22=2*1=2\]

OpenStudy (fifciol):

so you have \[\log_2x-2=5\] so \[\log_2x=7\]

OpenStudy (anonymous):

x = 2 x = 4 x = 32 x = 128

OpenStudy (fifciol):

\[\log_ab=c \rightarrow b=a^c\]

OpenStudy (fifciol):

try to write down these numbers in the form \[2^n\]

OpenStudy (fifciol):

for example

OpenStudy (fifciol):

\[2^5=32\]

OpenStudy (fifciol):

does it satisfy your equation?- no, because 5 is not equal to 7

OpenStudy (fifciol):

try to write down 128 in that form, can you do that?

OpenStudy (anonymous):

2^7

OpenStudy (fifciol):

yes

OpenStudy (fifciol):

log2x=7 does 2^7 satisfy this equation?

OpenStudy (fifciol):

using this formula you can check that: logab=c→b=a^c

OpenStudy (anonymous):

Great thanks! :)

OpenStudy (fifciol):

yw :)